# Towards Faithful and Robust LLM Specialists for Evidence-Based Question-Answering

### v3

@@ Appendix C Training Data Creation Process
| SynSciQA | 2143 | 6.11 | 112.08 | 21.64 | 4.28 | 27.22 | 1.93 | ||||||||||
| SynSciQA+ | 1386 | 6.27 | 112.21 | 21.57 | 4.20 | 26.78 | 1.65 | ||||||||||
| SynSciQA++ | 669 | 6.09 | 112.39 | 21.41 | 3.87 | 28.49 | 1.62 | ||||||||||
@@ Appendix D Hand-Evaluation of the Quality Filters
Appendix E Creation of GenSearch
@@ Appendix F Test Data Examples
To outline the differences between the test datasets, we further explore the properties of each evaluation benchmark.
The first evident difference lies in the statistical properties of the sources in the instructions (see Table 10). We create SynSciQA and GenSearch with the data creation pipeline described in Section 2.2. This means, we have known relevant and irrelevant sources in the datasets. On the other hand, we use top-10 retrieved sources for ChatReport and the output sources by ClimateQA for ClimateQA. As Table 10 shows, the different sourcing mechanisms result in a large difference in average length and standard deviation.
Furthermore, there are large differences in the structure and format of the sources. The exemplary comparison in Table 11 reveals that SynSciQA and GenSearch are predominantly in full-sentence form. While SynSciQA only contains synthetic scientific topics, GenSearch is created from internet sources and therefore much broader and more colloquial in tone. Sources in ChatReport start in the middle of the sentence, end in the middle of the sentence, and are not necessarily in full-sentence form. The same holds true for ClimateQA. However, this dataset also contains nested citations which represents the most complicated case for Evidence-Based QA.
@@
| SynSciQA | 539 | 6.10 | 98.23 | 21.28 | ||||
| GenSearch | 106 | 5.85 | 82.31 | 41.80 | ||||
| ChatReport | 110 | 10.00 | 67.94 | 25.27 | ||||
| ClimateQA | 261 | 4.47 | 134.51 | 13.35 | ||||
@@
| SynSciQA | The electromagnetic force is responsible for the most familiar interactions in our everyday lives. It is the force that allows us to see, feel, and interact with the world around us. When light interacts with matter, it can be absorbed, reflected, or transmitted, giving rise to the colors we perceive. The electromagnetic force also enables the operation of electronic devices, such as computers and smartphones, by allowing the flow of electric currents. Moreover, this force is essential for generating and transmitting electrical power. Understanding how particles interact through the electromagnetic force is not only of interest to scientists but also has practical applications that impact our daily lives. |
| GenSearch | But how?! Well, there’s merchandising, VOD, streaming video, foreign sales, and a plethora of other distribution channels that can help filmmakers, producers, and studios turn a profit. Traditionally, movies have made their money from ticket sales at the box office or in theaters. A studio might make about 60% of a film’s ticket sales in the United States, and around 20% to 40% of that on overseas ticket sales. |
| ChatReport | the details of Fortive’s climate-related governance, strategy, risk management, and metrics and targets. In 2021, we formally expanded the risk criteria within our Enterprise Risk Management program to account for the financial, operational, and regulatory risks in addition to physical risks for which we were already accounting. By incorporating additional climate-related risks into our existing protocol for evaluating and identifying risk, we are able to capture climate-related |
| ClimateQA | 5 that are subducted over decades are expected to experience significant warming (see Figure 5.3). The warming in the subtropical gyres penetrates deeper into the ocean than other gyres (roughly 15ºN–45ºN and 15ºS–45ºS in Figure 5.3), following the wind-driven bowing down of the density surfaces (the solid lines in Figure 5.3) in these gyres (Terada and Minobe, 2018). The greater warming at 700–2000 m in the Atlantic than the Pacific or Indian Oceans (Figure 5.3) reflects the strong southward transport of recently formed NADW at these depths by the AMOC. Two areas that commonly exhibit substantially reduced near-surface warming over the course of the 21st century are the northern north Atlantic, where a slowing AMOC (see Section 6.7.1.1) reduces the northward heat transport and brings the surface temperatures closer to what is found in other ocean basins at these latitudes (Collins et al. 2013), and the southern side of the Southern Ocean, where water upwells |
@@ Appendix G Hyperparameter and Other Settings
We always use random seed 42 for experiments in this work. We use the default QLoRA hyperparameter settings 555https://github.com/jondurbin/qlora, namely, an effective batch size of 32, a lora r of 64, a lora alpha of 16, a warmup ratio of 0.03, a constant learning rate scheduler, a learning rate of 0.0002, an Adam beta2 of 0.999, a max gradient norm of 0.3, a LoRA dropout of 0.1, 0 weight decay, a source max length of 2048, and a target max length of 512. We use LoRA module on all linear layers.
@@ Appendix H Relevance Label for Real RAG
Section 4 introduces that source-relevance label is available for SynSciQA thanks to the data creation process. We also annotate source-relevance for GenSearch. However, we do not annotate that for ChatReport and ClimateQA because we find most of the retrieved top-k sources in those real RAG systems are directly or indirectly relevant since they are retrieved from a narrow domain (e.g., a sustainability report). Take the following source-question pair as an example:
@@ Appendix I Statistical Significance Tests
To show the statistical significance of performance difference in Section 5.2, we first conduct Mann-Whitney U test on each sub-figure of Figure 4, Figure 5, Figure 6, and Figure 7. Specifically, we regard the scores of epoch 1 to 5 comes from the same distribution, and compute if distributions of different settings are statistically significantly different or not. For example, SynSciQA++ distribtuion ([81.56, 81.59, 80.83, 78.19, 81.9]) and SynSciQA distribtuion ([48.15, 62.01, 61.17, 57.05, 52.57]) in the first sub-figure of Figure 5. We use Mann-Whitney U test instead of student-t test to avoid making the normal distribution assumption. After having p-values between all settings, we apply Fisher’s method to aggregate the p-values, resulting in Table 2.
@@
GPT-3.5-turbo |
GPT-4 |
Llama-2-13b-chat |
Zephyr-7b- |
No Fine-Tuning |
Hum./GPT |
Hum./GPT |
Hum./GPT |
Hum./GPT |
| SynSciQA |
Hum./GPT |
Hum./GPT |
| SynSciQA |
GPT |
GPT |
| SynSciQA+ |
GPT |
GPT |
| SynSciQA+ |
GPT |
GPT |
| SynSciQA++ |
Hum./GPT |
Hum./GPT |
@@
The fine-tuning may only improve format quality as a short-cut to improving attributability scores, instead of making the answer sentences more citation-compliant. To verify this, we compute the attributability score again on those format-correct sentences only. In other words, all improvements should then be caused by more sentences supported by sources. The results are shown in Figure 11. It can be observed that in all settings, fine-tuning improves attributability without considering format quality. Interestingly, GPT-3.5 outperforms GPT-4 on ChatReport ignoring format quality, which coheres to findings in Ni et al. (2023) where GPT-3.5 is better entailed in ChatReport.

### v4

@@ Appendix C Training Data Creation Process
| SynSciQA | 2143 | 6.11 | 112.08 | 21.64 | 4.28 | 27.22 | 1.93 | ||||||||||||
| SynSciQA+ | 1386 | 6.27 | 112.21 | 21.57 | 4.20 | 26.78 | 1.65 | ||||||||||||
| SynSciQA++ | 669 | 6.09 | 112.39 | 21.41 | 3.87 | 28.49 | 1.62 | ||||||||||||
@@ Appendix D Hand-Evaluation of the Quality Filters
Appendix E Creation of GenSearchtest
@@ Appendix F Test Data Examples
To outline the differences between the test datasets, we further explore the properties of each evaluation benchmark. It is important to outline that we gradually leave distance from the properties of the in-domain dataset. This way, we ultimately aim to obtain insights into the real-world applicability of the approach.
The first evident difference lies in the statistical properties of the sources in the instructions (see Table 10). We create SynSciQAtest and GenSearchtest with the data creation pipeline described in Section 2.2. This means, we have known relevant and irrelevant sources in the datasets. On the other hand, we use top-10 retrieved sources for ChatReporttest and the output sources by ClimateQA for ClimateQAtest. As Table 10 shows, the different sourcing mechanisms result in a large difference in average length and standard deviation.
Furthermore, there are large differences in the structure and format of the sources. The exemplary comparison in Table 11 reveals that SynSciQAtest and GenSearchtest are predominantly in full-sentence form. While SynSciQAtest only contains synthetic scientific topics, GenSearchtest is created from internet sources and therefore much broader and more colloquial in tone. Sources in ChatReporttest start in the middle of the sentence, end in the middle of the sentence, and are not necessarily in full-sentence form. The same holds true for ClimateQAtest. However, this dataset also contains nested citations which represents the most complicated case for Evidence-Based QA.
@@
| SynSciQAtest | 539 | 6.10 | 98.23 | 21.28 | ||||
| GenSearchtest | 106 | 5.85 | 82.31 | 41.80 | ||||
| ChatReporttest | 110 | 10.00 | 67.94 | 25.27 | ||||
| ClimateQAtest | 261 | 4.47 | 134.51 | 13.35 | ||||
@@
| SynSciQAtest | The electromagnetic force is responsible for the most familiar interactions in our everyday lives. It is the force that allows us to see, feel, and interact with the world around us. When light interacts with matter, it can be absorbed, reflected, or transmitted, giving rise to the colors we perceive. The electromagnetic force also enables the operation of electronic devices, such as computers and smartphones, by allowing the flow of electric currents. Moreover, this force is essential for generating and transmitting electrical power. Understanding how particles interact through the electromagnetic force is not only of interest to scientists but also has practical applications that impact our daily lives. |
| GenSearchtest | But how?! Well, there’s merchandising, VOD, streaming video, foreign sales, and a plethora of other distribution channels that can help filmmakers, producers, and studios turn a profit. Traditionally, movies have made their money from ticket sales at the box office or in theaters. A studio might make about 60% of a film’s ticket sales in the United States, and around 20% to 40% of that on overseas ticket sales. |
| ChatReporttest | the details of Fortive’s climate-related governance, strategy, risk management, and metrics and targets. In 2021, we formally expanded the risk criteria within our Enterprise Risk Management program to account for the financial, operational, and regulatory risks in addition to physical risks for which we were already accounting. By incorporating additional climate-related risks into our existing protocol for evaluating and identifying risk, we are able to capture climate-related |
| ClimateQAtest | 5 that are subducted over decades are expected to experience significant warming (see Figure 5.3). The warming in the subtropical gyres penetrates deeper into the ocean than other gyres (roughly 15ºN–45ºN and 15ºS–45ºS in Figure 5.3), following the wind-driven bowing down of the density surfaces (the solid lines in Figure 5.3) in these gyres (Terada and Minobe, 2018). The greater warming at 700–2000 m in the Atlantic than the Pacific or Indian Oceans (Figure 5.3) reflects the strong southward transport of recently formed NADW at these depths by the AMOC. Two areas that commonly exhibit substantially reduced near-surface warming over the course of the 21st century are the northern north Atlantic, where a slowing AMOC (see Section 6.7.1.1) reduces the northward heat transport and brings the surface temperatures closer to what is found in other ocean basins at these latitudes (Collins et al. 2013), and the southern side of the Southern Ocean, where water upwells |
@@ Appendix G Hyperparameter and Other Settings
We always use random seed 42 for experiments in this work. We use the default QLoRA hyperparameter settings 666https://github.com/jondurbin/qlora, namely, an effective batch size of 32, a lora r of 64, a lora alpha of 16, a warmup ratio of 0.03, a constant learning rate scheduler, a learning rate of 0.0002, an Adam beta2 of 0.999, a max gradient norm of 0.3, a LoRA dropout of 0.1, 0 weight decay, a source max length of 2048, and a target max length of 512. We use LoRA module on all linear layers.
@@
This hyperparameter setup of training epochs orientates on previous impactful and practical work in the domain.777e.g., training for 2 epochs in https://aclanthology.org/2023.emnlp-main.245/ or for 3-5 epochs in https://github.com/tatsu-lab/stanford_alpaca or https://magazine.sebastianraschka.com/p/practical-tips-for-finetuning-llms that argue that multi-epoch training does not benefit LoRA. However, extending the study 5 to 10 or 15 epochs would likely make some arguments stronger.
@@ Appendix H Relevance Label for Real RAG
Section 4 introduces that source-relevance label is available for SynSciQAtest thanks to the data creation process. We also annotate source-relevance for GenSearchtest. However, we do not annotate that for ChatReporttest and ClimateQAtest because we find most of the retrieved top-k sources in those real RAG systems are directly or indirectly relevant since they are retrieved from a narrow domain (e.g., a sustainability report). Take the following source-question pair as an example:
@@ Appendix I Statistical Significance Tests
To show the statistical significance of performance difference in Section 5.2, we first conduct Mann-Whitney U test on each sub-figure of Figure 4, Figure 5, Figure 6, and Figure 7. Specifically, we regard the scores of epoch 1 to 5 comes from the same distribution, and compute if distributions of different settings are statistically significantly different or not. For example, SynSciQA++ distribtuion ([81.56, 81.59, 80.83, 78.19, 81.9]) and SynSciQAS distribtuion ([48.15, 62.01, 61.17, 57.05, 52.57]) in the first sub-figure of Figure 5. We use Mann-Whitney U test instead of student-t test to avoid making the normal distribution assumption. After having p-values between all settings, we apply Fisher’s method to aggregate the p-values, resulting in Table 2.
@@
| GPT-3.5-turbo | GPT-4 | Llama-2-13b-chat | Zephyr-7b- | |
| No Fine-Tuning | Hum./GPT | Hum./GPT | Hum./GPT | Hum./GPT |
| SynSciQA | - | - | Hum./GPT | Hum./GPT |
| SynSciQAS | - | - | GPT | GPT |
| SynSciQA+ | - | - | GPT | GPT |
| SynSciQA+S | - | - | GPT | GPT |
| SynSciQA++ | - | - | Hum./GPT | Hum./GPT |
@@
The fine-tuning may only improve format quality as a short-cut to improving attributability scores, instead of making the answer sentences more citation-compliant. To verify this, we compute the attributability score again on those format-correct sentences only. In other words, all improvements should then be caused by more sentences supported by sources. The results are shown in Figure 11. It can be observed that in all settings, fine-tuning improves attributability without considering format quality. Interestingly, GPT-3.5 outperforms GPT-4 on ChatReporttest ignoring format quality, which coheres to findings in Ni et al. (2023) where GPT-3.5 is better entailed in ChatReporttest. As a supplementary result, the improvement in format quality only is presented in Figure 12. Thus, the improvements in the Attributability score lie in both better formatting and entailment.